# Challenges and recommendations for Electronic Health Records data extraction and preparation for dynamic prediction modelling in hospitalized patients - a practical guide

### v1

@@ 4.1 Cohort definition
Defining the cohort of interest (e.g., hospital-wide population or patients with a specific condition) represents a first step in the prediction task definition. It can also be of interest during the project planning phase and it implies critical considerations (Table 4.4). An inaccurately defined cohort can lead to sample bias, resulting in performance estimates that do not accurately reflect the model’s performance in clinical practice.
@@ 4.2 Outcome definition
Prediction models using EHR data usually focus on in-hospital or post-discharge outcomes, including mortality, length of stay, readmission, acute events like bacteraemia, sepsis, and acute kidney injury, and chronic diseases such as heart failure, cancer, and cardiovascular disease (Pungitore and Subbian, 2023). We focus on outcomes that can be derived solely from structured EHR data, without linkage to external data sources or extraction of text notes, and are linked to a patient, i.e. we exclude resource utilization and workflow optimization outcomes (Table 4.4). Typically, a time window after the prediction trigger time is defined (also called prediction horizon) (Pungitore and Subbian, 2023), for both static (one prediction trigger) or dynamic (repeated prediction triggers throughout the patient’s stay) models.
@@ 4.3 Feature engineering
Feature engineering is the process of selecting, transforming, and creating features (variables) from the extracted EHR data to ensure that the data are brought into a suitable format for modelling and that relevant information is processed in a meaningful way. It includes data mapping and transformation (e.g.: grouping medical specialties in meaningful categories), converting timestamped events (e.g., lab results, medication administration, vital signs) into snapshot-based features and data aggregation (summarizing multiple observations or measurements into single features). The feature engineering is typically performed during data preparation, and most of times it happens concomitantly with data cleaning. Sometimes feature engineering is also performed at data extraction time, for example, for reducing the data volume for high-frequency time-series data, especially in ICU settings or for computing scores that are calculated and displayed in the EHR software but not stored in the system.
We distinguish between generic feature engineering (Table 4.4), which deals with the availability and mapping of clinical items, and time-sensitive feature engineering (Table 4.4), which deals with data aggregation for which correct processing of timestamps is critical and that can result in temporal leaks (i.e. data available at a specific time for training the model, but not available in the system at that timestamp). Time-sensitive feature engineering is critical for dynamic prediction models, can be useful also for some static models (e.g.: predictions at 24 hours after admission), and has less impact on models with prediction trigger at the end of the admission (e.g.: readmission prediction). We acknowledge that some minimal temporal leaks might not be very detrimental for the model performance or for its applicability, while others can have a large impact. However, following the DRIFT principle (Do It Right The First Time), good understanding of the extracted timestamps and correct handling of date/times during data extraction and preparation can safeguard against future problems, big or small.
@@ 4.4 Data cleaning
The data cleaning process constitutes of identifying and correcting data issues that could negatively impact the performance or applicability of prediction models. Errors in EHR data can arise from manual entry mistakes, improperly connected or malfunctioning devices, or bugs within the EHR software. Such errors are typically uncovered during data exploration and addressed during data preparation. Tools like the Data Quality Dashboard (Blacketer et al., 2021) have been developed to support and streamline the data cleaning process. Artifacts can also be introduced during data extraction or during data preparation. Unit tests for both the ETL process and the data preparation pipeline can safeguard against introducing errors.
@@
Overzealous correction of errors, such as manual correction during data preparation of every error encountered might not necessarily prove beneficial for prediction tasks, when such corrections cannot be programmatically reproduced during model implementation in clinical settings. This discrepancy can lead to a situation in which training data accurately represents the true patient state, while real-time predictions rely on erroneous data, resulting in reduced predictive performance. Although measures can be taken to address common and documented errors from previous studies or identified during data exploration (Table 4.4), it is impossible to anticipate and guard against all future errors. For example, new EHR software versions and updates may introduce new bugs, even as older bugs are resolved.
{longtabu}
to >p2cm>p5cm>p5cm>p1cm>p1cm COHORT Recommendations. ICU: Intensive Care Unit, EHR: Electronic Health Record, ICD: International Classification of Diseases, CPT: Current Procedural Terminology, PEEP: Positive End-Expiratory Pressure, FiO2: Fraction of Inspired Oxygen
Problem Problem description Recommendation METRIC cluster and dimension Weiskopf and Weng dimension
[…]
Problem Problem description Recommendation METRIC cluster and dimension Weiskopf and Weng dimension
[…]
Incorrect inclusion/exclusion criteria (and implicitly incorrect cohort definition) applied at extraction For some prediction tasks only a subset of hospitalised patients is of interest, e.g.: patients with a central catheter / urinary catheter, patients undergoing surgery (for postoperative outcomes), patients on mechanical ventilation, or patients with a specific diagnose or condition. The items that identify a device (a mechanical ventilator, a catheter) or the definitions used for defining a patient condition might differ between medical wards. Correct identification of these items at extraction time might be challenging and invertedly items may be missed, creating either gaps in data and incorrectly excluding admissions. Often it is difficult to spot such incorrect exclusions at data preparation time, as the researcher does not have access nor knowledge of the EHR system configuration, hospital workflow and data recording procedures in each department. Incorrect exclusion can be due to many factors, such as typos (an item has been configured with a typo for some period in the system, e.g.: centraal verneuze vs. veneuze katheter) or different recording procedures in specialized wards (e.g.: dialysis catheters are recorded in a special form in the Nephrology department), ultimately leading to sample selection bias (Sauer et al., 2022). If ethical approval can be granted to extract all admissions within the timeframe of interest, without applying other specific exclusion criteria, cohort definition can be done at data preparation time, as a collaborative effort between as data extraction engineers, researchers and clinical representatives. This allows researchers to perform additional verification, improving the cohort definition through multiple reviews. Whenever this is not possible, the data extraction engineers would have to work closely with clinical representatives for correct cohort definition. Correct clinical concept and terminology mapping becomes crucial. Representativeness - depth of data - coverage Completeness
Discontinuity or fragmentation of the patient stay due to cross-linking across different data sources Whenever the cohort definition is done at extraction, e.g.: admissions for patients with a specific device, and data are extracted from multiple databases (ICU and hospital-wide EHR), data for the entire hospital admission has to be extracted from both systems. For example, a patient admitted through emergency, transferred to ICU and then to Nephrology has a dialysis catheter only during the ICU stay. Extracting only the data during the ICU stay might not be sufficient for the prediction task. Data prior to ICU contains predictive information. Often, for devices like catheters, the patient is considered at risk 48 hours post catheter removal, therefore the patient is still at risk during the Nephrology stay. Alternatively, discontinuities can happen for extracted data when patient is transferred from ward to ward due to extraction of different items from more systems (e.g.: patient has a central catheter in internal medicine, is transferred to ICU, but a central line is not present, and gets transferred back to internal medicine ward, where the central catheter is present again) Whenever the hospital admission data is located in multiple data sources (ICU and EHR), data relating the entire hospital admission should be extracted, even if the inclusion criteria are met only in one of the data sources. Representativeness - depth of data - coverage Completeness
Inclusion/exclusion criteria require patient history prior to the study period Inclusion/exclusion criteria based on a history of chronic conditions, such as osteoarthritis (Thuraisingam et al., 2021), might not be fully recorded in the EHR system or would require extraction of data prior to the study period of interest, as historical conditions are linked to prior admissions. Assessing the completeness of data used to define the inclusion/exclusion criteria prior to study initiation can be done through external source linking to gold standard data sources (Thuraisingam et al., 2021). Alternatively, if historical diagnoses relevant to current admission are be available in (structured or unstructured) clinical notes recorded shortly after admission, these can be used. Representativeness - depth of data - coverage Completeness
Temporal leaks in cohort definition A subgroup of patients with a specific condition might be of interest for the prediction task, e.g.: cardiac patients, but predictions are made during the entire admission. Defining the subgroup using criteria that are not available at the start of admission (or at the first prediction trigger), such as discharge ICD codes (coded typically after patient discharge), or patient presence in the cardiology ward (while the patient was admitted through emergency, was transferred to ICU and then to cardiology), will affect the applicability of the model. The cohort can be defined using either data present at admission (e.g.: admission reason) or data available during previous admissions (e.g.: patients with ICD codes indicative of cardiovascular conditions during previous admissions). Alternatively, the first prediction trigger can be defined when data becomes available: predictions start at patient transfer to the cardiology department and end at patient discharge. Timeliness - currency Currency
Duplicated admissions present in extraction Due to extraction errors, the same patient admission is extracted twice, with different identifiers. This can occur due to extraction in separate batches or identifiers anonymization and leads to a biased cohort definition. Such problems can be captured in the data extraction framework using unit or integration tests. At data preparation, tables can be checked for identical duplicated records. Informativeness - redundancy - uniqueness NA
Restrictive inclusion/exclusion criteria applied at extraction limit the scope of missing data imputation Imputation of missing values (e.g.: lab results) might benefit from the extraction of a larger cohort than the one strictly of interest. Similarly, if the admissions are further split in episodes of interest for prediction, which are segments of time during a patient admissions when a patient is considered at risk for the outcome of interest (e.g.: catheter episodes, mechanical ventilator episodes), the imputation can be done in the largest possible context (using all patient-days during an admission instead of only the episode-patient-days). Performing missing data imputation in a larger context, would in theory, be beneficial, especially if the cohort of interest is rather small, although there is no evidence of this topic being studied. If ethical approval can be granted to extract all hospitalisations during a specified time frame, cohort definition using inclusion/exclusion criteria can be done at data preparation time, eventually after missing data imputation. Representativeness - depth of data - coverage Completeness
Restricting extraction to hospitalised patients Often, extraction is restricted to hospitalised patients (also called inpatient admissions), excluding outpatient (consultations) or day-care admissions. Some hospitalised patients will have data recorded during previous outpatient admissions. Depending on the hospital workflow, it is possible that at times (e.g.: for planned surgeries), pre-admission data is recorded on an outpatient admission that might be of clinical relevance for the current admission, e.g.: lab values might be ordered and linked to an outpatient admission that precedes the current inpatient admission with one or two days; these values might be considered recent enough to be useful for the prediction task. Extraction can be extended to all hospital admissions (inpatient, outpatient and day-care) and researchers can decide if outpatient or day-care admissions contain useful data for the prediction task. Representativeness - depth of data - coverage Completeness
@@
Linking identifiers As data sources are relational databases, the extraction process deals with a large number of identifiers. Patient, admission, physician and even bed (Maletzky et al., 2022) identifiers are used for linking clinical items. The process of anonymizing identifiers can also introduce errors. Furthermore, identifiers can differ in different systems (ICU vs. non-ICU). The process of denormalization at data extraction obscures a large number of identifiers. Nevertheless, generally patient and admission identifiers (anonymized) remain part of the extracted data. Incorrect linking of identifiers can results in incorrect extracted cohort. Correct linking of obscured identifiers can be tested at extraction time, for example by verifying date of birth of two linked patient items from two systems match (Ferrao et al., 2016). Representativeness - depth of data - coverage Completeness
@@
Definition of episodes of interest The patient follow-up time relevant for the prediction task might be different from the admission timeframe. This can either mean fragmenting hospital admissions in episodes of interest: catheter episode, mechanical ventilation episode, postoperative follow-up, or, conversely, merging subsequent overlapping hospital admissions (Miao et al., 2023). Defining the episodes of interest at data extraction time by using date filtering in large queries can prove error-prone Defining the episodes of interest during data preparation, rather than during data extraction, might prove less error-prone and comes with the advantage of benefiting from a larger patient context for prediction. Representativeness - depth of data - coverage Completeness
@@
Episode fragmentation Definition of episodes of interest, such as catheter episode and mechanical ventilator episodes can be done by start and end times: catheter placement and catheter removal, start and end time of mechanical ventilator, whenever these are available in the system and recorded at their time of occurrence. We warn that using the coded procedures (CPT - Current Procedural Terminology - codes) start and end times might hinder the model’s applicability, as these are generally used for billing and registered in the system post-factum, sometimes at the end of the admission. Whenever start and end times are not available, the registration of specific observations (e.g.: catheter monitoring observations) or parameters (PEEP and FiO2 for mechanical ventilation) identify an episode. These observations or parameters can exhibit gaps in registration. Observations can be recorded manually (catheters), or by devices (mechanical ventilation). Whenever recorded manually, they can display weekend effects. Exploration of the episodes can be done during data preparation, by checking the number of episodes per patient, the episodes length and time between episodes. Comparison to procedure coded placement and removal time can be done, whenever these are available, but we discourage using procedure codes in case they are not registered in the system in real time. Measurement process - completeness Completeness
{longtabu}
@@
to >p2cm>p5cm>p5cm>p1cm>p1cm OUTCOME Recommendations. ICD: International Classification of Diseases, COVID-19: Coronavirus Disease 2019, CAM: Confusion Assessment Method, DRS: Delirium Rating Scale, AKI: Acute Kidney Injury, CLABSI: Central Line-Associated Bloodstream Infection, VAE: Ventilator-Associated Event, CDC: Centers for Disease Control and Prevention, NA: Not Applicable.
Problem Problem description Recommendation METRIC cluster and dimension Weiskopf and Weng dimension
[…]
Problem Problem description Recommendation METRIC cluster and dimension Weiskopf and Weng dimension
[…]
@@
Outcomes derived based on ICD codes are not timestamped and often unreliable (under- or upcoded). Diagnoses are typically recorded by coders using ICD-9 or ICD-10 coding system, typically after patient discharge, without reflecting the time when the diagnose was established. In some cases through, these could be coded during the patient’s admission, although these situations are rare. Coded diagnoses are subject to undercoding, upcoding, or creep (misspecification or miscoding) (O’malley et al., 2005) or can represent the diagnostic differential rather than the actual diagnosis (Arbet et al., 2021; Ghassemi et al., 2020). Typically there will be more than one code that clinically identifies a disease (e.g.: chronic kidney disease) and these have to be carefully selected using clinical knowledge and mapped to the outcome label. Hospitals have also migrated from ICD9 to ICD10 historically, which might have introduced a period of discontinuity in the availability the diagnosis codes, and the coding practices might change after the transition to ICD-10 (O’malley et al., 2005). ICD codes might also evolve over time, e.g., introduction of a code for COVID-19 in 2020. Due to lack of timestamps, ICD codes are generally inappropriate for prediction of acute events during a patient’s admission. Differences in coding practices can introduce bias in the outcome defined by the ICD codes. When using ICD codes to define the outcome during a hospital admission (e.g.: acute events), it is important to understand the hospital’s coding practices, including potential under or upcoding and ensure that relevant timestamps for the event can be established. Additionally, it is important to explore temporal trends, as the coding practice might have changed over time. Measurement process - source credibility - traceability; Timeliness - currency Currency

### v2

@@ 4.1 Cohort definition
Defining the cohort of interest (e.g., hospital-wide population or patients with a specific condition) represents a first step in the prediction task definition (Table LABEL:tab:tbl-recommendations-cohort-1, Table LABEL:tab:tbl-recommendations-cohort-2 and Table LABEL:tab:tbl-recommendations-cohort-3). It can also be of interest during the project planning phase. An inaccurately defined cohort can lead to selection bias, resulting in performance estimates that do not accurately reflect the model’s performance in clinical practice.
@@ 4.2 Outcome definition
Prediction models using EHR data usually focus on in-hospital or post-discharge outcomes, including mortality, length of stay, readmission, acute events (like bacteremia, sepsis, and acute kidney injury), and chronic diseases (such as heart failure, cancer, and cardiovascular disease) (Pungitore and Subbian, 2023). We focus on outcomes that can be derived solely from structured EHR data, without linkage to external data sources or extraction of text notes, and are linked to a patient, i.e. we exclude resource utilization and workflow optimization outcomes (Table LABEL:tab:tbl-recommendations-outcome-1 and Table LABEL:tab:tbl-recommendations-outcome-2). Similar to cohort definition, a good practice represents assessing the feasibility of defining the outcome during the project planning phase.
@@ 4.3 Feature engineering
Feature engineering is the process of selecting, transforming, and creating features (variables) from the extracted EHR data to ensure that the data are brought into a suitable format for modelling and that relevant information is processed in a meaningful way. It includes data mapping (e.g., mapping medication brand name to generic drug names) and transformation (e.g., grouping medical specialties in meaningful categories), converting timestamped events (e.g., lab results, medication administration, vital signs) into snapshot-based features and data aggregation (summarizing multiple observations or measurements into single features). The feature engineering is typically performed during data preparation, and usually it happens concomitantly with data cleaning. Sometimes feature engineering is also performed at data extraction time, for example, for reducing the data volume for high-frequency time-series data, especially in ICU settings, or for computing scores that are calculated and displayed in the EHR software but not stored in the system. Clinical concept mapping, such as mapping medication to ATC (Anatomical Therapeutic Chemical) codes, laboratory tests to LOINC (Logical Observation Identifiers, Names, and Codes) codes or procedures and clinical observations to SNOMED CT (Systematized Nomenclature of Medicine Clinical Terms) can occur at different stages, including within the EHR software, during data extraction (e.g., OMOP CMD Standardized Vocabularies), or as part of data preparation.
We distinguish between generic feature engineering (Table LABEL:tab:tbl-recommendations-feat-not-ts-1 and Table LABEL:tab:tbl-recommendations-feat-not-ts-2), which deals with the availability and mapping of clinical items, and time-sensitive feature engineering (Table LABEL:tab:tbl-recommendations-feat-ts-1 and Table LABEL:tab:tbl-recommendations-feat-ts-2), which deals with data aggregation for which correct processing of timestamps is critical and that can result in temporal leaks (i.e. data available at a specific time for training the model, but not available in the system at that timestamp). Time-sensitive feature engineering is critical for dynamic prediction models, can be useful also for some static models (e.g., predictions at 24 hours after admission), and has less impact on models with prediction trigger at the end of the admission (e.g., readmission prediction). We acknowledge that some minimal temporal leaks might not be very detrimental for the model performance or for its applicability, while others can have a large impact. However, following the “do it right the first time” principle, good understanding of the extracted timestamps and correct handling of date/times during data extraction and preparation can safeguard against future problems, big or small.
As for the previous two groups, certain aspects of feature engineering can be evaluated during project planning, at least for the key features relevant to the prediction task and, at a minimum, for their availability. For instance, if a key feature was only recorded for a limited period before being discontinued or if timestamps of key features are unavailable for a dynamic prediction task, the project may become jeopardized.
@@ 4.4 Data cleaning
The data cleaning process constitutes of identifying and correcting data issues that could negatively impact the performance or applicability of prediction models. Errors in EHR data can arise from manual entry mistakes, improperly connected or malfunctioning devices, or bugs within the EHR software. Such errors are typically uncovered during data exploration and addressed during data preparation. Tools like the Data Quality Dashboard (Blacketer et al., 2021) have been developed to support and streamline the data cleaning process. Artifacts can also be introduced during data extraction or during data preparation. Unit tests for both the ETL process and the data preparation pipeline can safeguard against introducing additional errors.
@@
Overzealous correction of errors, such as manual correction during data preparation of every error encountered might not necessarily prove beneficial for prediction tasks, when such corrections cannot be programmatically reproduced during model implementation in clinical settings. This discrepancy can lead to a situation in which training data accurately represents the true patient state, while real-time predictions rely on erroneous data, resulting in reduced predictive performance. Although measures can be taken to address common and documented errors from previous studies or identified during data exploration (Table LABEL:tab:tbl-recommendations-clean-1 and Table LABEL:tab:tbl-recommendations-clean-2), it is impossible to anticipate and guard against all future errors. For example, new EHR software versions and updates may introduce new bugs, even as older bugs are resolved.
Table 1: COHORT Recommendations Related to Identifying Admissions. ICU: Intensive Care Unit; EHR: Electronic Health Record; HIS: Hospital Information System; METRIC = METRIC cluster and dimension mapping; $W$ = Weiskopf and Weng dimension mapping
Problem Description and recommendation
@@
Linking identifiers DESCRIPTION: As data sources are relational databases, the extraction process deals with a large number of identifiers. Patient, admission, physician and even bed (Maletzky et al., 2022) identifiers are used for linking clinical items. The process of anonymizing identifiers can also introduce errors. Furthermore, identifiers can differ in different systems (ICU vs. non-ICU). The process of denormalization at data extraction obscures a large number of identifiers. Nevertheless, generally patient and admission identifiers (anonymized) remain part of the extracted data. Incorrect linking of identifiers can result in incorrect extracted cohort.
RECOMMENDATION: Correct linking of obscured identifiers can be tested during data extraction, for example by verifying date of birth of two linked patient items from two systems match (Ferrao et al., 2016).
METRIC: Representativeness - depth of data - coverage; WW: Completeness
Cross-linking patients and admissions across various data sources creates inconsistencies DESCRIPTION: Patient demographics and admission data are typically stored in a HIS (Hospital Information System), which acts as a ’master’ system for patient data and shares these data with other specialized clinical software modules, such as the emergency system, ICU system, and EHR modules used in other wards. In a flawless integration between systems, all specialized clinical modules maintain identical copies of patient and admission information as the ’master’ HIS system. However, discrepancies might exist, especially in historical data. Issues such as failed patient merges, or demographic updates that do not correctly propagate across all systems, can lead to inconsistencies. Additionally, data migration from previous software vendors may introduce discrepancies. For example, a cohort including all patients in the emergency department over 60 might be differently identified based on HIS patient demographics than on the demographics from the emergency software. While these situations should normally be rare, they may be more common in historical data.
RECOMMENDATION: If large discrepancies have been created due to historical reasons, these would ideally be addressed during data extraction, considering the knowledge on historical system integration problems. If discrepancies persist in current clinical practice at the time of model deployment for clinical use, cohort identification and patient demographics may need to reflect the system in which the prediction model is used, e.g., if the model is deployed in the emergency department, patient data may need to reflect the information available to emergency physicians.
METRIC: Representativeness - depth of data - coverage; WW: Completeness
Invalid admissions, not reflecting real patients, are present in extraction DESCRIPTION: Admissions that do not represent real patients are present in EHR databases (de Kok et al., 2024). At minimum, ’test patients’ on which new EHR software versions are tested will be present in most EHR databases. Other examples can include: ’ghost’ admissions resulting from incorrect data migration from previous EHR vendor database, or administrative admissions for beds or equipment to enable lab sample collection. Such invalid admissions might or might not be linked to a patient record. Such admission can occur only during training data timeframe (for historical reasons, such as data migration problems), or both at training and clinical implementation time (for test patients).
RECOMMENDATION: Test patients can be easily filtered during data extraction by excluding a list of known test patient or admission identifiers; these can be difficult to detect during data preparation. Detecting invalid admissions due to historical problems requires knowledge of the database and the historical hospital workflows; generally, these are best handled during data extraction. If invalid admissions are detected during data exploration, the root cause of the underlying problem should better be discussed with data extraction engineers. Inclusion of such admissions at clinical implementation would generally not represent a real danger, as healthcare professionals are aware that predictions created for admissions that do not represent a real patient can be safely disregarded, and in most workflows, these would not interrupt their workflow. The impact on clinical implementation should though be evaluated case by case.
METRIC: Informativeness - redundancy - uniqueness; WW: NA
Duplicated admissions present in extraction DESCRIPTION: The same patient admission is extracted twice, with different identifiers. This can occur due to extraction in separate batches or identifiers anonymization and leads to a biased cohort definition. Duplicated admissions can also be recorded in the system for administrative or funding purposes.
RECOMMENDATION: Such problems can be captured in the data extraction framework using unit or integration tests. At data preparation, tables can be checked for identical duplicated records.
METRIC: Informativeness - redundancy - uniqueness; WW: NA
Restricting extraction to hospitalized patients may lead to omission of useful patient data DESCRIPTION: Often, data extraction is restricted to hospitalized patients (also called inpatient admissions), excluding outpatient (consultations) or day-care admissions. Some hospitalized patients may have relevant data recorded during previous outpatient admissions. Depending on hospital workflows, certain prediction tasks, such as those for planned surgeries, may benefit from pre-admission data recorded during outpatient visits. For instance, lab samples ordered one or two days before an inpatient admission may be linked to an outpatient visit and could be clinically relevant for the prediction task.
RECOMMENDATION: Extraction can be extended to all hospital admissions (inpatient, outpatient and day-care) and researchers can decide if outpatient or day-care admissions contain useful data for the prediction task.
METRIC: Representativeness - depth of data - coverage; WW: Completeness
Table 2: COHORT Recommendations Related to Inclusion/Exclusion Criteria. EHR: Electronic Health Record; ICD: International Classification of Diseases; METRIC = METRIC cluster and dimension mapping; $W$ = Weiskopf and Weng dimension mapping
Problem Description and recommendation
Incorrect or incomplete mappings leading to incorrect cohort definition applied at extraction DESCRIPTION: For some prediction tasks only a subset of hospitalized patients is of interest, e.g., patients with a central catheter / urinary catheter, patients undergoing surgery (for postoperative outcomes), patients on mechanical ventilation, or patients with a specific diagnose or condition. The items that identify a device (a mechanical ventilator, a catheter) or the definitions used for defining a patient condition might differ between medical wards. Correct identification of these items during data extraction might be challenging and inadvertently items may be missed, creating either gaps in data and incorrectly excluding admissions. Often it is difficult to spot such incorrect exclusions during data preparation, as the researcher does not have access nor knowledge of the EHR system configuration, hospital workflow and data recording procedures in each department. Incorrect exclusion can be due to many factors, such as typos (an item has been configured with a typo for some period in the system, e.g., centraal vs. central venous catheter) or different recording procedures in specialized wards (e.g., dialysis catheters are recorded in a special form in the Nephrology department), ultimately leading to sample selection bias (Sauer et al., 2022). For identifying patient conditions, multiple sources can be used, such as ICD codes or discharge diagnoses.
RECOMMENDATION: If ethical approval can be granted to extract all admissions within the timeframe of interest, without applying other specific exclusion criteria, cohort definition can be done during data preparation, as a collaborative effort between data extraction engineers, researchers and clinical representatives. This allows researchers to perform additional verification, improving the cohort definition through multiple reviews. Whenever this is not possible, the data extraction engineers would have to work closely with clinical representatives for correct cohort definition. Correct clinical concept and terminology mapping becomes crucial.
METRIC: Representativeness - depth of data - coverage; WW: Completeness
Inclusion/exclusion criteria require patient history prior to the study period DESCRIPTION: Inclusion/exclusion criteria based on a history of chronic conditions, such as osteoarthritis (Thuraisingam et al., 2021), may not be fully recorded in the EHR system or might require extracting data from a period preceding the study timeframe, for example, a ’look-back period’ of 3 to 5 years, as historical conditions are associated with previous admissions.
RECOMMENDATION: Assessing the completeness of data used to define the inclusion/exclusion criteria prior to study initiation can be done through external source linking to gold standard data sources (Thuraisingam et al., 2021). Alternatively, if historical diagnoses relevant to current admission are available in (structured or unstructured) clinical notes recorded shortly after admission, these can be used.
METRIC: Representativeness - depth of data - coverage; WW: Completeness
Restrictive inclusion/exclusion criteria applied at extraction limit the scope of missing data imputation DESCRIPTION: Imputation of missing values (e.g., lab results) might benefit from the extraction of a larger cohort than the one strictly of interest. Similarly, if the admissions are further split in episodes of interest for prediction, which are segments of time during a patient admissions when a patient is considered at risk for the outcome of interest (e.g., catheter episodes, mechanical ventilator episodes), the imputation can be done in the largest possible context (using all patient-days during an admission instead of only the episode-patient-days). Performing missing data imputation in a larger context, would in theory, be beneficial, especially if the cohort of interest is rather small, although there is no evidence of this topic being studied.
RECOMMENDATION: If ethical approval can be granted to extract all hospitalizations during a specified time frame, cohort definition using inclusion/exclusion criteria can be done during data preparation, eventually after missing data imputation.
METRIC: Representativeness - depth of data - coverage; WW: Completeness
Table 3: COHORT Recommendations Related to Episodes for Interest for Prediction. ICU: Intensive Care Unit, EHR: Electronic Health Record, ICD: International Classification of Diseases, CPT: Current Procedural Terminology, PEEP: Positive End-Expiratory Pressure, FiO2: Fraction of Inspired Oxygen; METRIC = METRIC cluster and dimension mapping; $W$ = Weiskopf and Weng dimension mapping
Problem Description and recommendation
Discontinuity or fragmentation of the patient stay due to cross-linking across different data sources DESCRIPTION: Whenever the cohort definition is done at extraction, e.g., admissions for patients with a specific device, and data are extracted from multiple databases (ICU and hospital-wide EHR), data for the entire hospital admission has to be extracted from both systems. For example, a patient admitted through emergency, transferred to ICU and then to the Nephrology ward has a dialysis catheter only during the ICU stay. Extracting only the data during the ICU stay might not be sufficient for the prediction task. Data prior to ICU contains predictive information. Often, for devices like catheters, the patient is considered at risk of catheter associated infections 48 hours post catheter removal, therefore the patient is still at risk for the event of interest for prediction post-ICU stay, during the stay at the Nephrology ward. Alternatively, discontinuities can happen for extracted data when patient is transferred between wards due to overlooked mapping of clinical items from more systems (e.g., virtually a patient has a central catheter in the general ward, is transferred to ICU without the presence of the central catheter, and gets transferred back to the general ward, with re-appearance of the central catheter).
RECOMMENDATION: Whenever the hospital admission data is located in multiple data sources (ICU and EHR), data relating the entire hospital admission should be extracted, even if the inclusion criteria are met only in one of the data sources.
METRIC: Representativeness - depth of data - coverage; WW: Completeness
Temporal leaks in cohort definition DESCRIPTION: A subgroup of patients with a specific condition might be of interest for the prediction task, e.g., cardiac patients, but predictions are made during the entire hospital admission. Defining the subgroup using criteria that are not available at the start of admission (or at the first prediction trigger), such as discharge ICD codes (coded typically after patient discharge), or patient presence in the cardiology ward (while the patient was admitted through emergency, was transferred to the ICU and then to the cardiology ward), will affect the applicability of the model.
RECOMMENDATION: The cohort should be defined using data available at first prediction trigger, e.g., data present at admission (e.g., admission reason) or data available during previous admissions (e.g., patients with ICD codes indicative of cardiovascular conditions during previous admissions). Alternatively, the first prediction trigger can be defined when data becomes available: predictions start at patient transfer to the cardiology department and end at patient discharge. Inclusion/exclusion criteria should never be defined using data after the prediction trigger, e.g., excluding patients who died for predictions during the patient’s hospitalization.
METRIC: Timeliness - currency; WW: Currency
@@
Definition of episodes of interest DESCRIPTION: The patient follow-up time relevant for the prediction task might be different from the admission timeframe, from admission time until discharge time. This can either mean fragmenting hospital admissions in episodes of interest: catheter episode, mechanical ventilation episode, postoperative follow-up, or, conversely, merging subsequent overlapping hospital admissions (Miao et al., 2023; de Kok et al., 2024). Defining the episodes of interest at data extraction time by using date filtering in large queries can prove error-prone.
RECOMMENDATION: Defining the episodes of interest during data preparation, rather than during data extraction, might prove less error-prone and comes with the advantage of benefiting from a larger patient context for prediction. Extraction of full admissions is recommended.
METRIC: Representativeness - depth of data - coverage; WW: Completeness
@@
Episode fragmentation DESCRIPTION: Definition of episodes of interest, such as catheter episode and mechanical ventilator episodes can be done by start and end times: catheter placement and catheter removal, start and end time of mechanical ventilator, whenever these are available in the system and recorded at their time of occurrence. We warn that using the coded procedures (CPT - Current Procedural Terminology - codes) start and end times might hinder the model’s applicability, as these are generally used for billing and registered in the system post-factum, sometimes at the end of the admission. Whenever start and end times are not available for prediction time, the registration of specific observations (e.g., catheter monitoring observations) or parameters (PEEP and FiO2 for mechanical ventilation) identify an episode. These observations or parameters can exhibit gaps in registration. Observations can be recorded manually in the EHR system (catheters), or recorded by devices and integrated in the EHR system with a prespecified granularity, like hourly or every minute (mechanical ventilation). Whenever recorded manually, they can display weekend effects. i.e. less or no values recorded during weekends due to understaffing and leading to data gaps. Whenever recorded by devices, the fragmentation might depend on the granularity of measurements integrated, which can be set differently per monitoring device.
RECOMMENDATION: Exploration of the episodes can be done during data preparation, by checking the number of episodes per patient, the episodes length and time between episodes. Comparison to procedure coded placement and removal time can be done, whenever these are available, but we discourage using procedure codes if these are not registered in the system in real time.
METRIC: Measurement process - completeness; WW: Completeness
@@
Table 4: OUTCOME Recommendations for Outcomes Directly Available in the Extraction. ICD: International Classification of Diseases; COVID-19: Coronavirus Disease 2019; CAM: Confusion Assessment Method; DRS: Delirium Rating Scale; AKI: Acute Kidney Injury; NA: Not Applicable; METRIC = METRIC cluster and dimension mapping; $W$ = Weiskopf and Weng dimension mapping
Problem Description and recommendation
@@
Outcomes derived based on ICD codes are not timestamped and often unreliable (under- or upcoded). DESCRIPTION: Diagnoses are typically recorded by coders using ICD-9 or ICD-10 coding system for reimbursing purposes, typically after patient discharge, without reflecting the time when the diagnose was established. In some cases through, these could be coded during the patient’s admission, although these situations are rare. Coded diagnoses are subject to undercoding, upcoding, or creep (misspecification or miscoding) (O’malley et al., 2005) or can represent the diagnostic differential (other possible conditions considered before confirming a final diagnosis) rather than the actual diagnosis (Arbet et al., 2021; Ghassemi et al., 2020). Coding practices can vary based on admission types; for example, ICD codes for emergency patients may be restricted only to the presenting complaint, while ICD codes for inpatients may reflect chronic diseases and comorbidities which are unrelated to the presenting complaint. Typically, there will be more than one code that clinically identifies a disease (e.g., chronic kidney disease) and these have to be carefully selected using clinical knowledge and mapped to the outcome label. Hospitals have also migrated from ICD9 to ICD10 historically, which might have introduced a period of discontinuity in the availability the diagnosis codes, and the coding practices might change after the transition to ICD-10 (O’malley et al., 2005). ICD codes might also evolve over time, e.g., introduction of a code for COVID-19 in 2020. Due to lack of timestamps, ICD codes are generally inappropriate for prediction of acute events during a patient’s admission. Differences in coding practices can introduce bias in the outcome defined by the ICD codes.
RECOMMENDATION: When using ICD codes to define the outcome during a hospital admission (e.g., acute events), it is important to understand the hospital’s coding practices, including potential under or upcoding and ensure that relevant timestamps for the event can be established. Additionally, it is important to explore temporal trends, as the coding practice might have changed over time.
METRIC: Measurement process - source credibility - traceability; Timeliness - currency; WW: Currency
Outcomes retrospectively assessed by clinical experts are labour-intensive and subject to human error. DESCRIPTION: Clinical experts can determine outcomes by reviewing patient records retrospectively, either systematically at regular intervals (e.g., an infection preventionist reviewing records monthly to report hospital-acquired infections to national public health institutes) or specifically for a prediction task (e.g., labeling patient records in extracted data). Both approaches require significant time and effort and are prone to errors.
RECOMMENDATION: One way to reduce human error is to involve a group of experts to reach a consensus on outcome assessment. However, this makes the task even more labour-intensive. Additionally, research or hospital teams may want to evaluate the model’s performance after clinical implementation, for example one year post-deployment. In this case, outcomes will need to be assessed by the same experts using the same procedure as during model training, which is rarely feasible.
METRIC: Measurement process - human-introduced error - noisy labels; Measurement process - completeness; WW: Correctness